OpenStudy (anonymous):

12 1/6 - 4 15/16 answer with mixed number if possible

OpenStudy (jdoe0001):

\({\bf 12\cfrac{1}{6}-4\cfrac{15}{16}\implies \cfrac{12\times 6+1}{6}-\cfrac{4\times 16+15}{16}=\cfrac{\qquad }{{\color{red}{ 48}}}\qquad } \begin{array}{llll} 48\div 6=8\\ 48\div 16=3\\ \hline\\ LCD={\color{red}{ 48}} \end{array}\)

OpenStudy (triciaal):

another approach separate whole numbers and fractions compute each then recombine and simplify 12-4 +1/6 - 15/16 LCD of 6 and 16 = 48 8 + (8+ 45)48 = 8 + 53/48 = 8 + 1+ 5/48 = 9 5/48